# Partially Stochastic Infinitely Deep Bayesian Neural Networks

### v2

@@ Abstract
In this paper, we present Partially Stochastic Infinitely Deep Bayesian Neural Networks, a novel family of architectures that integrates partial stochasticity into the framework of infinitely deep neural networks. Our new class of architectures is designed to improve the limitations of existing architectures around computational efficiency at training and inference time. To do this, we leverage the advantages of partial stochasticity in the infinite-depth limit which include the benefits of full stochasticity e.g. robustness, uncertainty quantification, and memory efficiency, whilst improving their limitations around computational complexity. We present a variety of architectural configurations, offering flexibility in network design including different methods for weight partition. We also provide mathematical guarantees on the expressivity of our models by establishing that our network family qualifies as Universal Conditional Distribution Approximators. Lastly, empirical evaluations across multiple tasks show that our proposed architectures achieve better downstream task performance and uncertainty quantification than their counterparts while being significantly more efficient.
@@ 1 Introduction
To overcome these challenges, previous research has explored partially stochastic networks as a potential solution in finite-depth settings (Daxberger et al. 2021b; Kristiadi et al. 2020; Li et al. 2024). These networks restrict stochasticity to a fraction of the model’s parameters, i.e. $\Theta=\Theta_{S}\cup\Theta_{D}$, where one learns both a distribution over $\Theta_{S}$ and point estimates for $\Theta_{D}$. These methods aim to achieve greater computational efficiency while retaining the benefits of Bayesian approaches. Initially, the main idea of these architectures was to approximate fully stochastic networks with a trade-off between computational efficiency and accuracy. However, Sharma et al. (2023) provide both theoretical and empirical evidence to demonstrate that partially stochastic networks can achieve comparable, if not superior, performance to their fully stochastic counterparts. While Sharma et al. (2023) leaves the selection of non-stochastic components in BNNs largely unexplored, works around Variational Bayesian Last Layers like Harrison et al. (2024), address this issue. As there is no trivial way to split the weights into subsets of deterministic and stochastic weights in the infinite-depth limit, previous research has only explored finite-depth regimes. However, studies such as Farquhar et al. (2020) hint towards the viability of extending this approach to an infinite-depth configuration.
Building on these insights, our work introduces Partially Stochastic Infinitely Deep Bayesian Neural Networks (PSDE-BNNs). This class of architectures combines the flexibility of neural differential equations with the efficient uncertainty quantification of partially stochastic networks. By selectively introducing noise within the weights process, our model achieves an improved representation of uncertainty and better performance in downstream tasks than its fully stochastic counterparts, while significantly reducing the computational overhead typically associated with infinitely deep BNNs. Our contributions to the field are summarised as follows:
@@ 2 Background and Preliminaries
where $\mu:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ (the drift), $\sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ (the volatility or diffusion matrix) are Lipschitz functions in the state, and satisfy a linear growth condition. Such conditions ensure the existence and uniqueness of a strong solution, given a square-integrable initial condition $X_{0}$. SDEs are a natural way of embedding stochasticity in a continuous-time framework, hence they are fit for modeling infinitely deep networks in a Bayesian setting. In infinite-depth Bayesian neural networks, Xu et al. (2022) propose to model the prior on weights $(w_{t})_{t\in[0,1]}$ as a SDE
@@
Variational inference on SDEs was introduced by Archambeau et al. (2007), and previous work have relied on other techniques to learn the coefficients of SDEs, such as Karhunen-Loève expansion of Brownian motion (Ghosh et al., 2022). The parametrization of the weights given by (2) require using KL divergence on path spaces. Minimizing the KL divergence between path measures induced by diffusion processes is a well-studied mathematical problem (Cattiaux & Léonard, 1994) which has been recently leveraged in the context of learning the Schrödinger Bridge Problem (Vargas et al., 2023) and regularizing generative models (Richter et al., 2024; Vargas et al., 2024). We use the following infinite-dimensional version of the KL divergence (Tzen & Raginsky, 2019; Li et al., 2020)
@@ 3 Method
where $t$ is the layer index, $h_{t}\in\mathbb{R}^{d_{h}}$ represents a vector of hidden unit activations at layer $t$, the input is $h_{0}=x$, $w_{t}$ are the parameters for layer t, and $\delta=1$. Taking $\delta=1/T$ and letting $T\rightarrow\infty$ yields a Neural ODE $\mathrm{d}h_{t}=f_{h}(t,h_{t};w_{t})\mathrm{d}t$, where $f_{h}$ is a neural network with parameters $w_{t}$. These parameters can also be modelled by a neural network (more precisely, a hypernetwork), which in the infinite-depth limit becomes itself a neural differential equation $\mathrm{d}w_{t}=f_{q}(t,w_{t};\theta)\mathrm{d}t$ (or, in the Bayesian setting, a Neural SDE). Therefore, $f_{q}$ is a neural network parametrised by $\theta$ which models the continuous evolution of the parameters of $f_{h}$. Finally, in this framework, the only trainable parameters are the vector $\theta$ and $w_{0}$. While we use SDE formalism similar to generative diffusion models, our model is discriminative. It ultimately learns the conditional distribution $p(y\mid x)$ via continuous latent variables.
@@ 3.1 Vertical separation of the weights
Notice that the diffusion of the approximate posterior is fixed to match the prior diffusion. This is to ensure that the KL divergence between the prior and posterior measures on paths $\mu_{p}$ and $\mu_{q}$ respectively, is finite. The proof of the following lemma can be found in (Archambeau et al., 2007).
@@
In conclusion. the dynamics of the whole system can then be written as a single SDE in $\mathbb{R}^{d_{h}+d_{w}}$:
@@ 4 Expressivity Guarantees
4.1 Constrained Stochastic Infinitely Deep Bayesian Neural Networks are not universal approximators
@@ 4.2 Partially Stochastic Infinitely Deep Bayesian Neural Networks are UCDAs
First, we assume that, for $m\in\mathbb{N}$ large enough, there exists a continuous generator function for the conditional distribution $Y|X$, i.e, a function $\tilde{f}:\mathcal{X}\times\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that for any $\eta\sim\mathcal{N}(0,I_{m})$ independent of $X$, $(X,\tilde{f}(X,\eta))$ is distributed like $(X,Y)$. This assumption is motivated by the so-called Noise Outsourcing lemma (Kallenberg, 2010) and we state it in Appendix A. Second, we use the results on the approximation capabilities of Neural ODEs, first proven by Zhang et al. (2020) - Lemmas 3 and 4 in Appendix A. Finally, we use the fact that classical Partially Stochastic finite-depth Bayesian Neural Networks are UCDAs (Sharma et al., 2023), as stated in Appendix A. We now present our main result:
@@ 5.1 Experimental Setup
We conduct our experiments under three distinct configurations. The first configuration, termed ‘SDEfirst’, is defined by setting the initial time cut, $t_{1}$, to 0. In this setup, the weights process is initially modelled with a SDE and then transitions to an ODE regime. Within this framework, we explore two sub-settings: one where the initial values of the ODE are the weights obtained from the SDE, and another, labelled ‘fix_w2’, where the initial values of the ODE are trainable parameters. The second configuration, ‘ODEFirst’, sets the second time cut, $t_{2}$, to 1, i.e. the endpoint of the considered time frame. Here, the weight evolution initially follows an ODE, with stochasticity introduced in the subsequent phase. The last configuration, the Horizontal Cut method (‘Hor. Cut’), splits the weight vector into two subsets, $S$ and $D$ here $t_{1}=0$ and $t_{2}=1$. These segments evolve according to an SDE, discretised using the Euler-Maruyama scheme, and an ODE computed using the Midpoint scheme, respectively. We emphasize two essential criteria on choosing $t_{1}$ and $t_{2}$. First is the data’s uncertainty and complexity. For data characterized by high uncertainty, or limited prior knowledge, extending the stochastic phase ($t_{2}$ in SDEFirst) may be beneficial. Conversely, when more prior information is available, prolonging the deterministic phase ($t_{1}$ in ODEFirst) could be more beneficial for the model. Second is the computational efficiency. Models employing a longer stochastic phase could demand greater computational resources, a factor that needs attention in scenarios with limited computational budgets. One can also refer to table 4 to see the impact of $t_{1}$ and $t_{2}$ on the performance and computing time of our models.
For hyperparameter optimisation, we employ a comprehensive grid search, which is detailed in Appendix E. Across all experiments, the diffusion function $g_{p}$ is set to a constant value. Lastly, experiments were conducted on a single Nvidia RTX 3090 GPU.
@@ 5.2 Image Classification
Similar to Xu et al. (2022), the hidden state’s instantaneous modifications ($f_{h}$) are parameterised with a convolutional neural network, which includes a strided convolution for downsampling and a transposed convolution layer for upsampling. The weights $w_{t}$ are set to be the filters and biases of these layers. The MLP parametrising $f_{q}$ includes bottleneck layers (e.g. 2-128-2) in order to minimise the variational parameters and ensure linear scalability with the dimension of $w_{t}$. For the MNIST (LeCun et al., 2010) dataset, a single SDE-BNN block is utilised, whereas, for CIFAR-10 (Krizhevsky, 2009), we employ a multi-scale approach with several SDE-BNN blocks with the invertible downsampling from Dinh et al. (2016) in between.
The classification outcomes, as detailed in Table 1, demonstrate the superior performance of our model against baseline models across the datasets. Notably, while ODEnet architectures (Hu et al., 2020) reach comparable classification metrics to standard residual networks, they fall short in terms of calibration. Our PSDE-BNN model exhibits enhanced accuracy relative to other stochastic frameworks, such as mean field variational inference (MFVI) ODEnets, which employ stochastic variational inference over weights invariant to depth. Furthermore, even in the realm of image classification, where discrete networks like ResNet32 with last-layer Laplace approximation exhibit strong performance, our continuous-depth approach remains competitive. It’s crucial to recognise the differences between these architectures; discrete networks may inherently align more closely with image classification tasks. However, our PSDE-BNN models achieve comparable performance without mechanisms like last-layer Laplace approximation, which are inapplicable in the infinite-depth setting. Moreover, the analysis reveals a consistent level of performance across the three distinct PSDE-BNN configurations, underlining the robustness of our approach regardless of the specific setting.
@@ 5.3 Uncertainty Quantification
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Table 1 shows that across the uncertainty-aware models, our PSDE-BNN (SDEFirst) surpasses all baseline models in ECE performance. While ODEFirst, SDEFirst with ‘fix_w2’, and PSDE-BNN (Hor. Cut) consistently demonstrate competitive error calibration and accuracy, SDE-BNN faces challenges in maintaining low ECE when striving for higher accuracy. This is because, during training, a trade-off emerges: ECE is sacrificed for improved accuracy, likely due to overfitting that narrows uncertainty estimates over time. The phenomenon is accentuated for the SDE-BNN which, at relatively low accuracies for CIFAR-10 (in the range of 70% - 80%) can achieve good ECE performance which quickly - and at a much faster rate than PSDE-BNNs - decays over epochs.
Encouraged by the calibration results, we now assess model performance in differentiating ID from OOD samples. A well-known problem with deterministic networks is that, regardless of the input sample, they will always be overconfident in their prediction. Ideally, this should not be the case for OOD samples, and models with good uncertainty quantification can reduce the magnitude of this issue. Our evaluation, based on ROC curves and AUC values (see Appendix B), shows that our partially stochastic models (AUC: 0.88 for models with vertical cuts, AUC:0.86 for the model with horizontal cut) outperform the SDE-BNN model (AUC: 0.84). Figure 3 further dissects the predictive entropy distributions, underscoring our models’ superior performance in representing OOD uncertainty.
Additionally, we have used CIFAR10-C (Hendrycks & Dietterich, 2019), a dataset which introduces various corruption levels to systematically assess our models’ uncertainty estimation. Our investigations included adding 17 different types of corruptions (e.g. gaussian noise, blur, snowification, compression, pixelation, etc) at 5 different levels of intensity to challenge the models’ robustness and generalisation. Results are shown in Table 2.
@@ 5.5 Ablation Studies
The ablation studies (Table 4 in Appendix C), show no correlation between the stochasticity ratio and downstream performance or uncertainty metrics (ECEs and entropy). This is consistent with the literature in partially stochastic networks, where several approaches achieve competitive uncertainty representations with only the last discrete layer of the network being stochastic (e.g. using Laplace approximation or HMC) demonstrating that a small subset of stochastic weights suffices (Daxberger et al. 2021a; Duran-Martin et al. 2022; Immer et al. 2021; Antorán et al. 2023). However, higher stochasticity ratios increase computation times. Given the lack of performance trade-off, the stochasticity ratio in our optimal model is set to 10%. Moreover, we find that increasing the order of $\kappa$ to $10^{-2}$ leads to an enhanced learning rate, but at the cost of numerical instability during training. In contrast, decreasing the order of $\kappa$ to $10^{-4}$ makes the training of the models converge slower. Given this trade-off, we conclude that the default value of $\kappa=10^{-3}$ is the most versatile (see Appendix C).

### v3

@@ Abstract
In this paper, we present Partially Stochastic Infinitely Deep Bayesian Neural Networks, a novel family of architectures that integrates partial stochasticity into the framework of infinitely deep neural networks. Our new class of architectures is designed to improve the computational efficiency of existing architectures at training and inference time. To do this, we leverage the advantages of partial stochasticity in the infinite-depth limit which include the benefits of full stochasticity e.g. robustness, uncertainty quantification, and memory efficiency, whilst improving their limitations around computational complexity. We present a variety of architectural configurations, offering flexibility in network design including different methods for weight partition. We also provide mathematical guarantees on the expressivity of our models by establishing that our network family qualifies as Universal Conditional Distribution Approximators. Lastly, empirical evaluations across multiple tasks show that our proposed architectures achieve better downstream task performance and uncertainty quantification than their counterparts while being significantly more efficient. The code can be found at https://github.com/Sergio20f/part_stoch_inf_deep
@@ 1 Introduction
To overcome these challenges, previous research has explored partially stochastic networks as a potential solution in finite-depth settings (Daxberger et al. 2021b; Kristiadi et al. 2020; Li et al. 2024). These networks restrict stochasticity to a subset of the model’s parameters, denoted as $\Theta=\Theta_{S}\cup\Theta_{D}$, and the learning process involves estimating a distribution for $\Theta_{S}$ while obtaining point estimates for $\Theta_{D}$. These methods aim to achieve greater computational efficiency while retaining the benefits of Bayesian approaches. Initially, the main idea of these architectures was to approximate fully stochastic networks with a trade-off between computational efficiency and accuracy. However, Sharma et al. (2023) provide both theoretical and empirical evidence to demonstrate that partially stochastic networks can achieve comparable, if not superior, performance to their fully stochastic counterparts. While Sharma et al. (2023) leaves the selection of non-stochastic components in BNNs largely unexplored, works focusing on Variational Bayesian Last Layers like Harrison et al. (2024), address this issue. As there is no trivial way to split the weights into subsets of deterministic and stochastic weights in the infinite-depth limit, previous research has only explored finite-depth regimes. However, studies such as Farquhar et al. (2020) hint towards the viability of extending this approach to an infinite-depth configuration.
Building on these insights, our work introduces Partially Stochastic Infinitely Deep Bayesian Neural Networks (PSDE-BNNs). This class of architectures combines the flexibility of neural differential equations with the efficient uncertainty quantification of partially stochastic networks. By selectively introducing noise into the weights process, our model achieves an enhanced representation of uncertainty and superior performance in downstream tasks compared to fully stochastic counterparts. Additionally, it significantly reduces the computational overhead typically associated with infinitely deep BNNs. Our contributions to the field are summarised as follows:
@@ 2 Background and Preliminaries
where $\mu:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ (the drift), $\sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ (the volatility or diffusion matrix) are Lipschitz functions in the state, and satisfy a linear growth condition. Such conditions ensure the existence and uniqueness of a strong solution, given a square-integrable initial condition $X_{0}$. SDEs are a natural way of embedding stochasticity in a continuous-time framework, hence they are fit for modeling infinitely deep networks in a Bayesian setting. For these networks, Xu et al. (2022) propose to model the prior on weights $(w_{t})_{t\in[0,1]}$ as an SDE
@@
Variational inference on SDEs was introduced by Archambeau et al. (2007), and previous work has relied on other techniques to learn the coefficients of SDEs, such as the Karhunen-Loève expansion of Brownian motion (Ghosh et al., 2022). The parametrization of the weights given by (2) requires using KL divergence on path spaces. Minimizing this divergence between path measures induced by diffusion processes is a well-studied mathematical problem (Cattiaux & Léonard, 1994) which has been recently leveraged in the context of learning the Schrödinger Bridge Problem (Vargas et al., 2023) and regularizing generative models (Richter et al., 2024; Vargas et al., 2024). We use the following infinite-dimensional version of the KL divergence (Tzen & Raginsky, 2019; Li et al., 2020)
@@ 3 Method
where $t$ is the layer index, $h_{t}\in\mathbb{R}^{d_{h}}$ represents a vector of hidden unit activations at layer $t$, the input is $h_{0}=x$, $w_{t}$ are the parameters for layer t, and $\delta=1$. Taking $\delta=1/T$ and letting $T\rightarrow\infty$ yields a Neural ODE $\mathrm{d}h_{t}=f_{h}(t,h_{t};w_{t})\mathrm{d}t$, where $f_{h}$ is a neural network with parameters $w_{t}$. These parameters can also be modelled by a neural network (more precisely, a hypernetwork), which in the infinite-depth limit becomes itself a neural differential equation $\mathrm{d}w_{t}=f_{q}(t,w_{t};\theta)\mathrm{d}t$ (or, in the Bayesian setting, a Neural SDE). Therefore, $f_{q}$ is a neural network parametrised by $\theta$ which models the continuous evolution of the parameters of $f_{h}$. Finally, in this framework, the only trainable parameters are the vector $\theta$ and $w_{0}$. While we use an SDE formalism similar to generative diffusion models, our model is discriminative. It ultimately learns the conditional distribution $p(y\mid x)$ via continuous latent variables.
@@ 3.1 Vertical Separation of the Weights
Notice that the diffusion of the approximate posterior is fixed to match the prior diffusion. This is to ensure that the KL divergence between the prior and posterior measures on paths $\mu_{p}$ and $\mu_{q}$ respectively, is finite. This can be formulated in the following lemma (see Archambeau et al., 2007 for a proof).
@@
In conclusion, the dynamics of the whole system can then be written as a single SDE in $\mathbb{R}^{d_{h}+d_{w}}$:
@@ 4 Expressivity Guarantees
4.1 Constrained Infinitely Deep Bayesian Neural Networks are not Universal Approximators
@@ 4.2 Partially Stochastic Infinitely Deep Bayesian Neural Networks are UCDAs
First, we assume that, for $m\in\mathbb{N}$ large enough, there exists a continuous generator function for the conditional distribution $Y|X$, i.e, a function $\tilde{f}:\mathcal{X}\times\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that for any $\eta\sim\mathcal{N}(0,I_{m})$ independent of $X$, $(X,\tilde{f}(X,\eta))$ is distributed like $(X,Y)$. This assumption is motivated by the Noise Outsourcing lemma (Kallenberg, 2010) and we state it in Appendix A. Second, we use the results on the approximation capabilities of Neural ODEs, first proven by Zhang et al. (2020) - Lemmas 3 and 4 in Appendix A. Finally, we use the fact that Partially Stochastic finite-depth Bayesian Neural Networks are UCDAs (Sharma et al., 2023), as stated in Appendix A. We now present our main result:
@@ 5.1 Experimental Setup
We conduct our experiments under three distinct configurations. The first configuration, termed ‘SDEfirst’, is defined by setting the initial time cut, $t_{1}$, to 0. In this setup, the weights process is initially modelled with an SDE and then transitions to an ODE regime. Within this framework, we explore two sub-settings: one where the initial values of the ODE are the weights obtained from the SDE, and another, labelled ‘fix_w2’, where the initial values of the ODE are trainable parameters. The second configuration, ‘ODEFirst’, sets the second time cut, $t_{2}$, to 1, i.e. the endpoint of the considered time frame. Here, the weight evolution initially follows an ODE, with stochasticity introduced in the subsequent phase. The last configuration, the Horizontal Cut method (‘Hor. Cut’), splits the weight vector into two subsets, $S$ and $D$ here $t_{1}=0$ and $t_{2}=1$. These segments evolve according to an SDE, discretised using the Euler-Maruyama scheme, and an ODE computed using the Euler scheme. Two critical criteria guide the selection of $t_{1}$ and $t_{2}$. The first criterion is the data’s uncertainty and complexity. For data with high uncertainty or limited prior knowledge, extending the stochastic phase ( $t_{2}$ in ‘SDEFirst’) may be advantageous. Conversely, when more prior information is available, prolonging the deterministic phase ( $t_{1}$ in ‘ODEFirst’) could benefit the model more. The second criterion is computational efficiency. Models with an extended stochastic phase may require greater computational resources, which is an important consideration in scenarios with limited computational budgets. Table 4 provides further insights into the impact of $t_{1}$ and $t_{2}$ on the performance and computational time of our models.
We employ a comprehensive grid search for hyperparameter optimisation, and the final values are detailed in Appendix E. Across all experiments, the diffusion function $g_{p}$ is set to a constant value. Lastly, experiments were conducted on a single Nvidia RTX 3090 GPU.
@@ 5.2 Image Classification
Similar to Xu et al. (2022), the hidden state’s instantaneous modifications ($f_{h}$) are parameterised with a convolutional neural network, which includes a strided convolution for downsampling and a transposed convolution layer for upsampling. The parameters $w_{t}$ are set to be the weights and biases of these layers. The MLP parametrising $f_{q}$ includes bottleneck layers (e.g. 2-128-2) in order to minimise the variational parameters and ensure linear scalability with the dimension of $w_{t}$. For the MNIST (LeCun et al., 2010) dataset, a single SDE-BNN block is utilised, whereas, for CIFAR-10 (Krizhevsky, 2009), we employ a multi-scale approach with several SDE-BNN blocks with the invertible downsampling from Dinh et al. (2016) in between.
The classification outcomes, as detailed in Table 1, demonstrate the superior performance of our model against baseline models across the datasets. Notably, while ODEnet architectures (Hu et al., 2020) achieve classification metrics comparable to standard residual networks, they exhibit inferior calibration. In contrast, our PSDE-BNN model demonstrates superior accuracy relative to other stochastic frameworks, such as mean field variational inference (MFVI) ODENets, which utilize stochastic variational inference (SVI) over depth-invariant weights. Moreover, even in the domain of image classification, where discrete networks like ResNet32 with last-layer Laplace approximation perform robustly, our continuous-depth approach remains competitive. It’s crucial to recognise the differences between these architectures; discrete networks may inherently align more closely with image classification tasks. However, our PSDE-BNN models achieve comparable performance without mechanisms like last-layer Laplace approximation, which are inapplicable in the infinite-depth setting. Moreover, the analysis reveals a consistent level of performance across the three distinct PSDE-BNN configurations, underlining the robustness of our approach regardless of the specific setting.
@@ 5.3 Uncertainty Quantification
Table 1 demonstrates that among the uncertainty-aware models, our PSDE-BNN (SDEFirst) achieves superior Expected Calibration Error (ECE) performance compared to all baseline models. Although ODEFirst, SDEFirst with ‘fix_w2’, and PSDE-BNN (Hor. Cut) consistently exhibit competitive error calibration and accuracy, the SDE-BNN encounters difficulties in maintaining low ECE when pursuing higher accuracy. This trade-off arises during training, where ECE is compromised for enhanced accuracy, likely due to overfitting that progressively narrows uncertainty estimates. This phenomenon is particularly pronounced in the SDE-BNN. For CIFAR-10, at relatively low accuracies (70% - 80%), SDE-BNN can achieve good ECE performance, but this quickly degrades over epochs, and at a much faster rate than PSDE-BNNs. This is shown in the last row of Table 1, where SDE-BNN, within the same time limit, achieves excellent ECE at the cost of reduced accuracy, which would improve with additional epochs.
Encouraged by the calibration results, we proceed to evaluate model performance in distinguishing in-distribution (ID) from out-of-distribution (OOD) samples. A well-known issue with deterministic networks is their tendency to be overly confident in their predictions, regardless of the input. Ideally, models with effective uncertainty quantification should mitigate this problem for OOD samples. Our evaluation, based on ROC curves and AUC values (see Appendix B), demonstrates that our partially stochastic models (AUC: 0.88 for models with vertical cuts, AUC: 0.86 for the model with horizontal cut) outperform the SDE-BNN model (AUC: 0.84). Figure 3 further analyzes predictive entropy distributions, highlighting our models’ superior performance in representing OOD uncertainty.
Moreover, we utilized CIFAR10-C (Hendrycks & Dietterich, 2019), a dataset that introduces various corruption levels to systematically assess our models’ uncertainty estimation. Our investigations involved adding 17 different types of corruptions (e.g., Gaussian noise, blur, snow, compression, pixelation) at five different intensity levels to challenge the models’ robustness and generalization. The results are presented in Table 2.
@@ 5.5 Ablation Studies
The ablation studies (Table 4 in Appendix C), show no correlation between the stochasticity ratio and downstream performance or uncertainty metrics (ECEs and entropy). This is consistent with the literature in partially stochastic networks, where several approaches achieve competitive uncertainty representations with only the last discrete layer of the network being stochastic (e.g. using Laplace approximation or HMC) demonstrating that a small subset of stochastic weights suffices (Daxberger et al. 2021a; Duran-Martin et al. 2022; Immer et al. 2021; Antorán et al. 2023). However, higher stochasticity ratios increase computation times. Given the lack of performance trade-off, the stochasticity ratio in our optimal model is set to 10%. Moreover, we find that increasing the order of $\kappa$ to $10^{-2}$ leads to an enhanced learning rate but at the cost of numerical instability during training. In contrast, decreasing the order of $\kappa$ to $10^{-4}$ makes the training of the models converge slower. Given this trade-off, we conclude that the default value of $\kappa=10^{-3}$ is the most versatile (see Appendix C).